Here's the problem:
The thickness of the wing of an adult pixie is 4.5 x 10^-2. The wing of a newborn pixie is 1.5 x 10^-5 inches thick. How many times thicker is the is the adults wing than the newborns?
How much larger than the adults?
Is the thickness solution this?
4.5 x 10^-2 / 1.5 x 10^-5
4.5/1.5 x 1/10 ^-7
3 x 10^-7
how do I find how much larger is the adults?

Found 2 solutions by MathLover1, Alan3354:
Answer by MathLover1(20849)
Answer by Alan3354(69443)

Here's the problem:
The thickness of the wing of an adult pixie is 4.5 x 10^-2. The wing of a newborn pixie is 1.5 x 10^-5 inches thick. How many times thicker is the is the adults wing than the newborns?
How much larger than the adults?
Is the thickness solution this?
4.5 x 10^-2 / 1.5 x 10^-5
4.5/1.5 x 1/10 ^-7
3 x 10^-7
how do I find how much larger is the adults?
===============
Or, you can do it like this:
4.5 x 10^-2 / 1.5 x 10^-5 = 0.045/0.000015 = 3000 times AS LARGE
= 2999 times larger.
------------------
I would do the the way the other tutor did, tho.
Especially if the exponents are greater, eg, 6.28*10^23, or 3*10^8.